OpenStudy (anonymous):

i still need help...we had a substitute today and didn't cover the math homework :( my question is there are a total of 20 territories, denmark has 2, and norway has 1/5 of uk. so i figured out the answer is uk=15 and norway =3 because 3=1/5 of 15 and 15+3=18 but i don't know how to set this up as an equation and would like to know the proper way of doing it can you help me please :)

OpenStudy (anonymous):

let's say that the amount of territories of the uk is x then the amount of territories of norway is 0.2x we know that denmark has 2, and that there are 20 total. so the uk and norway combined is 18. so x + 0.2x = 1.2x = 18 x = 18/1.2 = 15 so the uk has 15 territories and norway had 0.2 * 15 = 3 territorries last check: 15 (uk) + 3 (norway) + 2 (denmark) = 20, as it should be.

OpenStudy (anonymous):

thank you so much, it looks easy now that you explained it :)